OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 8?

OpenStudy (amistre64):

do you recall the general formula for a circle?

OpenStudy (amistre64):

it is just the pythag thrm using different letters x^2 + y^2 = r^2

OpenStudy (amistre64):

since r = radius; and r^2 = 8 ....

OpenStudy (anonymous):

?

OpenStudy (anonymous):

okayy nevermind i got it